[ Three times the least of three consecutive odd INTEGERS exceeds twice the greatest by 17. Find the three Integers. ]
Answer by TimothyLamb(4379)

3i = 2k + 17
k = i + 4
j = i + 2
---
3i = 2k + 17
3i = 2(i + 4) + 17
3i = 2i + 8 + 17
---
i = 25
j = 27
k = 29
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php